OpenStudy (anonymous):

medals a jewelry salesman earns a commission of 25% of all jewelry sales made.Yesterday a soul five bracelets for $60 each and four rings for $40 each how much did he earn in commission yesterday

OpenStudy (texaschic101):

5 bracelets for $60 each......5 * 60 = 300 4 rings for $40 each...4 * 40 = 160 for a total of : (300 + 160) = 460 the salesman earns 25% commission off of all sales... (turn percent to decimal) 0.25 * 460 = $ 115 So he earned $115 off of his $460 sales yesterday.